One-third(d+9)=-12 what is d??
Answer by 303795(602)

The general procedure is to isolate the d. Undo all the operations that have happened to the d in the reverse order. For this problem d is added to the 9 first (because it is in brackets) then it was multiplied by one third. The order in which to undo the operations will be undo times by one third first then undo adding to 9.

Multiply both sides of the equation by 3 first (undoes the multiply by one third)
3 * One-third(d+9)=3 * (-12)
d + 9 = -36 Subtract 9 from each side (undoes the add 9)
d + 9 - 9 = -36 - 9
d = -45

Check that one third of (-45 + 9) = -12